Please help me solve this problem. I would appreciate step by step so that it can help me with other exercises.
-6x+2x+6+5x=/0-9/-/-6+5/

Answer by gonzo(654)

i'm assuming those are parentheses rather than slashes.
your equation looks to me like it should be:
-6x + 2x + 6 + 5x = (0-9) - (-6+5)
if you really meant slashes then i have no idea what you are talking about.
assuming you meant parentheses, this would be solved as follows:
-----
first you solve within the parentheses going from left to right using the hierarchy of algebraic operations (exponentiation, multiplication and division, addition and subtraction).
-----
since 0-9 = -9, and -6 + 5 = -1, the equation becomes:
-6x + 2x + 6 + 5x = (-9) - (-1)
-----
you combine like terms by performing indicated algebraic operations on them.
-----
-6x + 2x + 5x becomes 7x - 6x becomes x
-9 - (-1) becomes -9 + 1 becomes -8
-----
after you have combined like terms on each side of the equation, it becomes:
x + 6 = -8
-----
you subtract 6 from both sides of the equation to get
x = -8 - 6
which becomes
x = -14
-----
you prove your solution is correct by substituting -14 for x in the original equation.
-----
-6x + 2x + 6 + 5x = (0-9) - (-6+5) becomes
(-6*-14) + (2*-14) + 6 + (5*-14) = (0-9) - (-6+5)
performing all the indicated operations within the parentheses makes the equation become:
84 - 28 + 6 - 70 = -9 - (-1)
combining like terms on each side of the equation by performing all the indicated operations on them makes the equation become:
-8 = -9+1
which becomes
-8 = -8
since both sides of the equation are the same, the value of x = -14, when substituted in the equation, proves to be correct.
-----
gonzo@gmx.us